From the edge of the roof of a building that is 30 m high, a stone was thrown horizontally 

with a velocity of 15 m/s. How long did it take the stone to touch the ground?


1
Expert's answer
2021-06-13T07:09:00-0400
"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}},""t=\sqrt{\dfrac{2\cdot30\ m}{9.8\ \dfrac{m}{s^2}}}=2.47\ s."